A restaurant receives bread from two bakeries, A and B. Bakery A is on time sixty five percent of the time. Bakery B is on time fifty two percent of the time. What is the probability that one bakery will be late
Answer by robertb(5830)

P(A' or B') = P(A') + P(B') - P(A' and B')
=0.35 + 0.48 - 0.35*0.48 = 0.662